OpenStudy (anonymous):

Find the value of y so that so that the line passing through (4, y) and (5, 3) has a slope of 1. A. 6 B. 1 C. -2 D. 2

OpenStudy (across):

Solve for \(y\):\[1=\frac{3-y}{5-4}\]

OpenStudy (anonymous):

so... 3y-1?

OpenStudy (across):

No.\[1=\frac{3-y}{5-4}\\1=3-y\\y=2\]

OpenStudy (anonymous):

ohh, alright, thanks